Shift summary for Day, Dec. 30, 1999

Sasha Glazov (glazov@hep.uchicago.edu)
Thu, 30 Dec 1999 15:25:18 -0600

Shift Summary for day shift, 30/Dec/99

On Shift: VOD,SG

Hours of beam delivered at all intensities: 4.1
Hours of beam delivered at nominal intensity: 3.9
Hours of E799 data (special runs not included): 3.7
Hours of special runs (laser scans, muon runs...): 0.

Accounting of lost time: MI Power problem
Run stop/start because of L3 accept error

========================================================================
Runs ended this shift:
Run 15321 -- 5.0M
Run 15322 -- 6.3M

Tapes brought to FCC this shift:
none

========================================================================
Accesses during shift: few, after the end of 99 running.

========================================================================
Overview of shift:
E799 datataking after until 12.00
Preparation for Y2K/startup in January afterwords

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15321 102 7:55 Sydney's voice asking us to check L3 accept fraction
It looks bad for ktev4. Stop the run, fresh start --
15322 1 8:07 Start new run
9:12 MI power suply problems 10min estimate
9:45 New estimate: 45 min
79 9:52 Beam is back !!! Low intensity though.
97 10:15 Back to the nominal intensity
12:00 Last proton in 1999.

>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>

Now a lot of people in the control room doing differnet things.
Sydney&VOD start system backup. Sydney flashes cachers.
Leo&SashaL swaped ADC for later KAMI use in crates 11&12.
Run pedestal run without problems. Got permit lost in Cosmic configuration
Took crates 11/12 out of readout --> still have problems.
Finally it starts working but with beam triggers 1 & 8 only (but this
is enough for JasonL).

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Do not forget to put ADC crates 11 & 12 back into readout